Osage, Inc., manufactures and sells lamps. The company produces only when it receives orders and, therefore, has no inventories. The following information is available for the current month:

  


Actual (based on actual orders for 457,000 units)
Master Budget (based on budgeted orders for 494,000 units)
Sales revenue
$4,975,000



$4,940,000

Less










Variable costs










Materials

1,475,000




1,482,000

Direct labor

283,000




345,800

Variable overhead

675,100




642,200

Variable marketing and administrative

482,000




494,000

Total variable costs
$2,915,100



$2,964,000

Contribution margin
$2,059,900



$1,976,000

Less










Fixed costs










Manufacturing overhead

990,200




960,700

Marketing

295,000




295,000

Administrative

211,000




180,700

Total fixed costs
$1,496,200



$1,436,400

Operating profits
$563,700



$539,600


 

Required:

Prepare a profit variance analysis for Osage, Inc. (Do not round intermediate calculations. Indicate the effect of each variance by selecting "F" for favorable, or "U" for unfavorable. If there is no effect, do not select either option.)
Answer #1

flexible budget column

sales revenue 457,000(from actual question)*(4,940,000(master budget column)/ 494,000(master budget question)= 4,570,000 

variable costs: 

materials 457,000*(1,482,000/ 494,000)= 1,371,000

direct labor 457,000*(345,800/ 494,000)= 319,900 

Variable overhead: 457,000*(642,200/ 494,000)= 594,100 

Variable marketing and administrative: 457,000*(494,000/ 494,000)= 457,000 

Total variable costs 457,000*(2,964,000/ 494,000)= 2,742,000

Contribution margin 457,000*(1,976,000/ 494,000)= 1,828,000 

Fixed costs:

Manufacturing overhead 960,700

Marketing 295,000

Administrative 180,700

Total fixed costs 960,700+ 295,000+ 180,700= 1,436,400

Operating profits 1,828,000- 1,436,400= 391,600


sales activity variable column

flexible budget - master budget, except all fixed cost section


sales price variance

sales revenue 4,975,000(actual)- 4,570,000(flexible)= 405,000 F

Cm= 405,000 F

operating profits= 405,000 F


market and administrative variance

variable costs:

Variable marketing and administrative 482,000(actual)- 457,000(flexible)= 25,000u

total variable cost  25,000 U

CM= 25,000 U

Fixed Costs:

administrative 211,000- 180,700= 30,300 U

total fixed cost= 30,300 U

operating profits 25,000+ 30,300= 55,300 U


Manufacturing variances

variable costs:

material 1,475,000(actual)- 1,371,000(flexible)= 104,000 U

direct labor= 283,000- 319,900= 36,900 F

variable overhead= 675,100- 594,100= 81,000 U

total variable cost= 104,000- 36,900+ 81,000= 148,100 U

CM= 148,100 U

fixed costs:

manufacturing overhead 990,200- 960,700= 29,500 U

total fixed cost= 29,500 U

operating profits 148,100+ 29,500= 177,600